# It’s Morphing Time: Unleashing the Potential of Multiple LLMs via Multi-objective Optimization

### v1

@@ 4.2.2 Case Study
Table 2 provides a comparative analysis of responses to a mathematical reasoning problem by three models, illustrating the remarkable improvement in performance through the introduction of a multi-objective optimization algorithm in our MM-MO method. By employing multi-objective optimization methods to search for the optimal configuration for model merging, we not only enhanced our model’s performance in the multi-disciplinary evaluation test (C-EVAL, Table 1) and code generation (Human Eval, Table 1) but also further uncovered its mathematical reasoning capabilities.
@@
4.2.3 Ablation Study
Table 3: Ablation study: (1) DARE-TIES without MM-MO (2) MM-MO, consider only the accuracy of C-EVAL (3) MM-MO, consider both the accuracies of C-EVAL and GSM8K (Default setting).
Merging Methods Source Models C-EVAL GSM8K Human Eval
DARE + TIES Single Model 1 + 2 + 3 70.0 51.32 50.60
MM-MO
(1 obj: C-EVAL) Single Model 1 + 2 + 3 71.5 66.18 51.82
MM-MO
(2 objs: C-EVAL & GSM8K) Single Model 1 + 2 + 3 71.4 66.56 56.09
In this ablation study, we compared the impact of different model merging methods on model performance, specifically comparing DARE + TIES and two types of MM-MO (single-objective and multi-objective) methods. The table presents the performance of these three merging methods across three evaluation tasks (C-EVAL, GSM8K, and Human Eval).
First, using the default configuration (average weights), the DARE + TIES method serves as the baseline, achieving scores of 70.0, 51.32, and 50.60 on C-EVAL, GSM8K, and Human Eval, respectively. Next, we introduced the proposed MM-MO method. Initially, we considered the MM-MO method with a single objective focused solely on C-EVAL. This method increased the score on C-EVAL to 71.5, indicating that our black-box optimization strategy in model merging can enhance performance on specific tasks. Moreover, since C-EVAL is a comprehensive academic assessment test set, improving the accuracy of the merged model on C-EVAL also brought enhancements in mathematical reasoning and code generation capabilities. Specifically, the model obtained through the single-objective MM-MO merging strategy achieved scores of 66.18 and 51.82 on GSM8K and Human Eval, respectively, outperforming our baseline.
Finally, we employed the MM-MO method considering both C-EVAL and GSM8K as objectives. This approach yielded scores of 71.4, 66.56, and 56.09 on C-EVAL, GSM8K, and Human Eval, respectively. Compared to the single-objective MM-MO method focusing only on C-EVAL, this approach achieved comparable results on C-EVAL and a further notable improvement in the accuracy on GSM8K. This infers that the mathematical reasoning (GSM8K) capability of LLM may contribute to code generation (Human Eval) ability. Additionally, although Human Eval was not the best, the increased number of optimization objectives enhanced the model’s overall capabilities, significantly improving its code generation ability, with the score increased to 56.09. This demonstrates that considering multiple objectives can achieve better trade-off and overall performance across different tasks.
In conclusion, the proposed MM-MO method, by incorporating multi-objective black-box optimization, effectively enhances model performance across various tasks, underscoring the importance and advantages of multi-objective optimization in model merging configurations.
5 Limitation
Our experiments were constrained by the available computational resources, limiting the application of our method to the merging of models with up to 7 billion parameters. While our approach demonstrated promising results within this parameter range, the effectiveness and scalability of our method for models with significantly larger parameter counts remain untested. Future research is necessary to evaluate the performance and feasibility of our multi-objective optimization method for merging models with higher parameter volumes. This exploration will be crucial to understand the full potential and limitations of our approach in handling more extensive and complex models.
Additionally, our model merging method currently requires that the source models be homologous, meaning they must be fine-tuned versions of the same pre-trained model. This homogeneity constraint ensures compatibility during the merging process. However, it also limits the applicability of our method to scenarios where source models are derived from different pre-trained models. Future work should explore the feasibility of extending our approach to handle heterogeneous source models, broadening the scope and utility of our model merging technique.
6 Conclusion
This paper proposes a novel approach for large language model merging using multi-objective black-box optimization algorithms, addressing two major challenges: the reliance on human intuition for merging configurations and the potential loss of crucial parameters with DARE strategies. By employing qEHVI, our method automates the search for optimal merging configurations, ensuring that the merged model maintains high performance without the typical parameter conflict issues.
Model merging offers significant advantages over training models from scratch, including enhanced knowledge integration, performance optimization, and cost efficiency. Our approach effectively combines specialized knowledge from multiple pre-trained models, creating a unified model that leverages the best qualities of each, resulting in superior overall performance. Additionally, merging models is more resource-efficient, requiring less energy and computational power.
Our contributions include automating the model merging process, enhancing model potential by preserving individual capabilities and improving performance on diverse tasks, and validating our method through extensive evaluations on various natural language tasks. The results demonstrate that our approach achieves superior performance compared to existing methods, offering a robust, high-performing solution for integrating diverse task-specific models.
In summary, our multi-objective optimization method provides an automated, efficient, and powerful solution for model merging, addressing key challenges and resulting in robust, high-performing models that leverage the strengths of multiple pre-trained models.

### v2

@@ C.2 Case Study: Mathematical Reasoning
Table 6 provides a comparative analysis of responses to a mathematical reasoning problem by three models. By employing MM-MO to search for the optimal configuration for model merging, we enhanced our model’s performance in mathematical reasoning capabilities.
@@
C.3 Parameter Sensitivity Experiment
Table 7: Performance of MM-MO with different number of function evaluations.
Merging Method Models Average Score C-EVAL C-EVAL (hard) GSM8K HumanEval
Single Model 1 Qwen1.5-7B-Chat 55.34 68.7 51.1 54.59 46.95
Single Model 2 Liberated-Qwen1.5-7B 55.47 69.7 50.1 53.30 48.78
Single Model 3 firefly-qwen1.5-en-7b 51.01 70.0 50.7 49.81 33.54
MM-MO / FE = 20 Single Model 1 + 2 + 3 54.62 70.7 51.5 53.60 42.68
MM-MO / FE = 40 Single Model 1 + 2 + 3 59.27 72.0 52.8 57.62 54.66
MM-MO / FE = 30 (Default) Single Model 1 + 2 + 3 59.47 71.9 52.7 57.77 55.49
In this section, we compare the performance of MM-MO with different number of function evaluations to validate the sensitivity of our proposed method to this parameter. Table 7 illustrates the impact of varying number of function evaluations on the performance of MM-MO.
When FE=20, the performance of the merged model across various tasks is consistently inferior to that of FE=30 and FE=40. This suggests that a function evaluation count of 20 is insufficient to discover an optimal model merging configuration. Comparing the performances of the merged model between FE=30 and FE=40, we observe that while the accuracy on the C-EVAL under FE=30 is lower than under FE=40, its performance on GSM8K and HumanEval slightly surpasses that of the latter. This indicates that a function evaluation count of 30 is sufficient for MM-MO to search for a satisfactory model merging configuration. Therefore, considering the trade-off between computational cost and performance, we set the default number of function evaluations in this study to 30.
C.4 Time Overhead of MM-MO
Table 8: Time overhead of MM-MO for different model size.
Merging Method Model Size Time Overhead (min)
MM-MO 1.5B 259.21
MM-MO 7B 526.52
Table 8 lists the time overhead of MM-MO for different model size, 1.5B and 7B. Compared to fine-tuning on large datasets to improve the overall performance of the model, using MM-MO for model merging with existing models does not require additional training data. It only needs to be evaluated on the existing language model evaluation datasets to quickly obtain a merged model with great overall performance in a relatively short period of time.
C.5 The Final Configuration Searched using MM-MO
Figure 4: Final Configuration Searched using MM-MO. (1.5B & 7B)
Figure 4 illustrates the final parameter configuration of model merging (1.5B & 7B) searched using MM-MO. Density is the fraction of weights in differences from the base model to retain and weight represents the relative weighting of a given model. From the figure we can know the contribution of different single models to the final model merging.
Appendix D Licenses
The licenses for the codes we used in this work are shown in Table 9.
Table 9: List of licenses for the codes suite we used in this work.
Resources Type Link License
MergeKit [42] Code https://github.com/arcee-ai/mergekit LGPL-3.0 license
OpenCompass [43] Code https://github.com/open-compass/opencompass Apache-2.0 license